# Cyber Food Swamps: Investigating the Impacts of Online-to-Offline Food Delivery Platforms on Healthy Food Choices

### v2

@@ Abstract
Online-to-offline (O2O) food delivery platforms have substantially enriched the food choices of urban residents by allowing them to conveniently access farther food outlets. However, concerns about the healthiness of delivered food persist, especially because the impact of O2O food delivery platforms on users’ healthy food choices remains unclear. This study leverages large-scale empirical data from a leading O2O delivery platform to comprehensively analyze online food choice behaviors and how they are influenced by the online exposure to fast food restaurants, i.e., online food environment. Our analyses reveal significant discrepancy in food preferences across demographic groups and city sizes, where male, low-income, and younger users and those located in larger cities more likely to order fast food via O2O platforms. Besides, we also perform a comparative analysis on the food exposure differences in online and offline environments, confirming that the extended service ranges of O2O platforms can create larger “cyber food swamps”. Furthermore, regression analysis highlights that a higher ratio of fast food orders is associated with “cyber food swamps”, areas characterized by a higher share of accessible fast food restaurants. A 10% increase in this share raises the probability of ordering fast food by 22.0%. Moreover, a quasi-natural experiment substantiates the long-term causal effect of online food environment changes on healthy food choices. Our findings underscore the need for O2O food delivery platforms to address the health implications of online food choice exposure, thereby informing efforts by various stakeholders to improve residents’ dietary health.
@@ Introduction
Although the convenience of O2O food delivery is undeniable, concerns persist regarding the nutritional quality of delivered food, with numerous criticisms pointing to its generally low nutritional value (Stephens, Miller, and Militello 2020; Dai, Wu, and Hu 2022). Drawing on the concept of ’food swamps,’ which describes areas oversaturated with unhealthy food outlets in the physical world (Stowers, Schwartz, and Brownell 2017), O2O food delivery platforms have the potential to expand food access and create virtual food swamps that shape consumer choices towards unhealthy restaurants, leading to adverse health outcomes (Stephens, Miller, and Militello 2020). Therefore, it is essential to quantitatively assess how online food environments on these platforms influence users’ health food choices.
Existing studies primarily used surveys and field studies to evaluate O2O food delivery platforms from the perspective of healthiness, including the nutritional value of delivered food (Brar and Minaker 2021; Dai, Wu, and Hu 2022), consumer perceptions of healthy food availability (Dai, Wu, and Hu 2022; Eu and Sameeha 2021; Osaili et al. 2023), and healthy food choices (Osaili et al. 2023; Giacomini et al. 2024). These studies often suffered from in the scale of research samples and lack a direct link between online food environment and consumers’ real-world food choices. Furthermore, while significant efforts have been made to examine the impact of traditional dining environments on food choices (Feng et al. 2010; Althoff et al. 2022; García Bulle Bueno et al. 2024), their findings may not directly correspond to the effects of online food environments. This is particularly relevant as O2O food delivery significantly expands the spatial range of food options, highlighting the need to compare online and offline food swamps.
@@
To answer the above questions, we conduct comprehensive data-driven analyses and derive insightful findings based on O2O food delivery restaurant and consumption data at both city and individual levels, collected from one of the largest O2O food delivery platforms in China. First, we focus on the macro-level health discrepancies within the O2O food delivery industry. We uncover a scaling law that governs the relation between O2O food delivery usages and city sizes, showing larger cities have significantly higher proportions of fast food restaurants and placed delivery orders, which may be attributed to the faster pace of social life in these cities (Bettencourt et al. 2007). Healthiness of food choices is also correlated with demographic, with males, individuals with lower incomes, and younger people exhibiting a significantly higher likelihood of ordering fast food online. This trend, derived from large-scale empirical data, mirrors previous findings on food destinations based on small-sample surveys (Kerr et al. 2012). Second, we compare the distributional patterns between offline and online food environments. The number of accessible O2O restaurants is significantly higher than that of local offline restaurants, while the healthiness of the online food environment is slightly better than that of the offline environment. However, the extended service range of O2O platforms expands the spatial coverage of food swamps. The online food environment shows a stronger correlation with healthy food choices on O2O food delivery platforms. Finally, following the analytic framework proposed by García Bulle Bueno et al. (García Bulle Bueno et al. 2024), we examine the relationship between the online food environment and healthy food choices. Logistic regressions reveal that a higher proportion of fast food restaurants in a user’s location is associated with a higher likelihood of ordering fast food. Notably, males, low-income individuals, and younger users are more affected by this environment, underscoring the heterogeneous susceptibility of users on online platforms (Sukiennik, Gao, and Li 2024; Li et al. 2022). A quasi-natural experiment based on users who permanently change their ordering context further substantiates the gradual, negative causal impact of online food environments dominated by unhealthy fast food options on healthy dining habits, highlighting the urgent need for regulations addressing “cyber food swamps”.
Our main contributions can be summarized as follows:
@@
1. We leverage empirical O2O food delivery consumption data to comprehensively understand the healthiness of the online food environment and its impact on healthy food choices, highlighting the significant potential for exploring the interplay between the web and society using data-driven methods.
@@
2. We demonstrate macro-level health discrepancy within the online food environment by uncovering city-level scaling laws of O2O food delivery adoption and identifying disparities in healthy food choices across gender, age, and income groups.
@@
3. We compare the healthiness of offline and online food environments and confirm the amplifying effect of O2O food delivery on real-world food swamps. 4. We reveal the effects of online food environments on healthy food choice behaviors through an quasi-intervention experiment utilizing a detailed, individual order-level dataset, indicating the negative health impacts of “cyber food swamps”.
@@
Many previous studies and reports have substantiated the rapid growth of O2O food delivery in global cities, highlighting it as a convenient and viable solution to address limited access to healthy food options (Dillahunt, Simioni, and Xu 2019; Saarijärvi, Mitronen, and Yrjölä 2014; Yeo, Goh, and Rezaei 2017). During the COVID-19 pandemic, social distancing measures made O2O food delivery a primary means of obtaining food, further accelerating the industry’s growth (Dsouza and Sharma 2021; Li, Mirosa, and Bremer 2020; Hong et al. 2021). At the same time, the nutritional and health implications of O2O food delivery have attracted increasing attention from researchers worldwide. As for users’ perceptions of delivered food, surveys conducted in Jordan (Osaili et al. 2023), Malaysia (Eu and Sameeha 2021), and China (Dai, Wu, and Hu 2022) found that most participants believe food delivered through O2O food delivery is generally less healthy than food served in restaurants. However, many of these participants still report a high reliance on such services. Several studies have evaluated the healthiness of O2O food delivery meals through sampling. An analysis of food delivery menus in Canada revealed that the majority of menus scored below the standards set by the Healthy Eating Index (Brar and Minaker 2021). Similarly, a study on the nutritional quality of food sold on Chinese O2O food delivery platforms found that most popular items have low nutritional value. Younger consumers, in particular, tend to pay less attention to the nutritional content of delivered meals, which can lead to health issues such as elevated cholesterol and obesity with long-term consumption (Dai, Wu, and Hu 2022). In Italy, an analysis of differences in willingness to use food delivery services among users found that individuals with lower health literacy are more dependent on O2O food delivery (Giacomini et al. 2024). Most current studies are limited to small-scale data collection through surveys or sampling, without assessing the overall health preferences of O2O food delivery users across a broader population. In our study, we leverage a large-scale empirical dataset covering the consumption behaviors of hundreds of thousands of users to comprehensively analyze their online food environments and healthy food choices.
@@
Digital data collected from online services, including activity tracking, GPS traces, and social media posts, have been widely linked to the health outcomes of urban residents in existing research. For instance, large-scale physical activity data collected via smartphone applications reveal the relationship between discrepancy in daily step counts and obesity indicators across various countries. This discrepancy significantly decreases as city walkability improves (Althoff et al. 2017). A study based on mobile GPS data found that while large cities offer more recreational and social options, they tend to direct individuals to locations frequented by people of similar socioeconomic status, increasing the segregation of social activities and leading to new public health risks, such as lack of physical activity and imbalanced diets (Nilforoshan et al. 2023; Chen et al. 2023). Social media data, spanning multiple modalities, are also emerging as promising proxies for public health indicators. Posts on X (formerly Twitter) have been used to predict county-level public health metrics such as obesity and diabetes rates (Culotta 2014; Abbar, Mejova, and Weber 2015). Food-related photos on Instagram, combined with location information, can accurately reflect offline food deserts (De Choudhury, Sharma, and Kiciman 2016). YouTube videos have been applied to track dietary changes during the COVID-19 pandemic (Mejova and Manikonda 2023). Additionally, online recipe platforms are being used to nudge healthier food adoption by promoting specific healthy recipes (Chelmis and Gergin 2023; Jesse, Jannach, and Gula 2021). Similarly, data generated by O2O food delivery platforms hold the potential to reflect the health-related information of urban residents.
@@
In this study, we mainly use following three categories of data, i.e., O2O food delivery restaurant data, O2O food delivery consumption order data, and demographic data.
@@
We collect information on O2O food delivery restaurants from one of China’s largest O2O food delivery platforms. For each restaurant, the platform provides its name, category, and service range. To evaluate the healthiness of the restaurants, we classify those within the “fast food and snacks”, “western fast food”, and “fried food” categories, as these are often regarded as having low nutritional value and may contribute to health issues such as obesity (Hu et al. 2016). All other categories are considered non-fast food restaurants. On average, 18.77% of restaurants are categorized as fast food restaurant.
@@
The service range of each O2O food delivery restaurant indicates the urban areas that can access its delivery services. As depicted in Figure 1, the restaurant marked in blue has a broader service range than the one marked in pink. Utilizing this data, we can identify O2O food delivery restaurants accessible to each urban area. We divide the city of Beijing into Geohash-6 “regions” (approximately 1.2$\times$0.6 kilometers), represented as black rectangles in Figure 2. For each region $c$, the healthiness of its corresponding online food environment $\phi(c)$ is defined as the ratio of the number of accessible fast food restaurants to the total number of accessible restaurants. Figure 2(a) illustrates the spatial distribution of $\phi(c)$ across 4,921 regions in Beijing.
@@
The city-level dataset record the number of daily O2O food delivery orders, including fast food orders, across 283 Chinese cities. For the individual-level dataset, we randomly sample 850,000 platform users in Beijing and collect their order records from January 2017 to May 2018. Each entry includes the user ID, order time, user location (at the Geohash-6 region level), expenditure amount, and the corresponding O2O food delivery restaurant. Each user placed 50.3 orders on average. Based on these detailed order records, we calculate the fast food preference $\psi$ of each user, defined as the ratio of fast food orders to the total number of orders. Figure 2(b) illustrates the distribution of $\psi$ on the sampled population. On average, 27.98% of all O2O food delivery orders are placed at fast food restaurants.
@@
To establish a connection between healthy food choices and user attributes, we further obtain demographic data from the platform, which includes gender, income level, and age, as inferred by the platform’s machine learning algorithms from user behaviors. To ensure user privacy, income levels are categorized into three groups: “high-income”, “medium-income” and “low-income”. Similarly, age groups are segmented into “below 25”, “25 to 40”, and “over 40”. The demographic distributions of the sampled population is depicted in Figure 3.
@@
We first demonstrate the macroscopic city-level health discrepancy of O2O food delivery. Urban services and economic activities often follow a scaling law, expressed as $Y_{i}=Y_{0}N_{i}^{\alpha}$ (Bettencourt 2021; Ribeiro and Rybski 2023), where $Y_{i}$ represents a measure of socioeconomic activity or resources, and $N_{i}$ denotes the city’s population. An exponent $\alpha<1$ indicates sublinear growth of $Y$ relative to the city population, typically associated with urban infrastructure such as gas stations or road length, where the per-capita infrastructure decreases in larger cities (Um et al. 2009). In contrast, socioeconomic factors like GDP or crime rates often exhibit $\alpha>1$, signifying superlinear growth, where per-capita values increase more rapidly in larger cities due to intensified social interactios (Bettencourt 2013; Succar and Porfiri 2024).
Figure 4: Scaling laws of O2O food delivery restaurants (a), fast food restaurants (b), O2O food delivery orders (c), and fast food orders (d) in Chinese cities.
We collect city population from China City Statistical Yearbook [^1] https://www.stats.gov.cn/sj/ndsj/2019/indexch.htm and analyze the scaling law of O2O food delivery. As an innovative form of urban services, O2O food delivery in Chinese cities also exhibits scaling behavior. The fitted lines in Figure 4(a) and (c) indicate that the scaling exponent $\alpha$ for the number of O2O food delivery restaurants and the number of O2O food delivery orders are 0.86$\pm$0.07 and 1.17$\pm$0.10, respectively. This corresponds to the typical sublinear and superlinear growth patterns of urban infrastructure and economic activities, with $\beta=\frac{5}{6}$ and $\frac{7}{6}$, reflecting higher infrastructure utilization efficiency in large cities (Bettencourt 2013).
Furthermore, the scaling exponents for fast food restaurants and orders are significantly higher than those for all food categories ($P<0.05$), as shown in Figure 4(b) and (d). Specifically, doubling the city population increases the share of fast food restaurants by 4.0% and the share of fast food orders by 5.4%. This difference suggests a health discrepancy across cities, with larger cities demonstrating a greater reliance on fast food restaurants through O2O food delivery. This may be due to the high efficiency of fast food chains in fast-paced urban environments (Bettencourt et al. 2007).
@@
The observed health discrepancies across cities of different sizes may be related to variations in the composition of urban populations. To test this assumption, we further investigate the relationships between user demographics and their behaviors on O2O food delivery platforms. We focus on three demographic attributes - gender, income level, and age—and compare the average proportion of fast food orders and the average order price for each group. The values and corresponding 95% confidence intervals are shown in Figure 5.
@@
Figure 5: The average fast food order ratio and order price across different gender (a), income level (b), and age (c) groups. Error bars represent 95% confidence intervals.
@@
The proportion of fast food orders among male users (33.0%) is significantly higher than that of female users (25.0%), while the average price of food delivery orders is nearly identical for males (58.4 CNY) and females (58.6 CNY). This indicates that male users have a noticeably stronger preference for fast food delivery compared to female users. In terms of income level, the average order price increases with higher income, while the proportion of fast food orders decreases. Users of highest income level only have 26.0% fast food orders, while the ratio for lowest income level reaches 32.0%. This suggests that higher-income groups tend to opt for more expensive, healthy food delivery options, whereas lower-income groups prefer cheaper, less healthy food options. Similarly, the average price of food delivery orders rises with age, while the proportion of fast food orders declines from 34.8% among users under 25 years to 24.8% for users over 40 years, indicating that younger people are more inclined to choose fast food, while older individuals tend to prefer higher-priced alternatives.
@@
This overall health discrepancy highlights the heterogeneous healthy food choices of O2O food delivery platform users, which may be influenced by factors such as service pricing (Li et al. 2024), the online food environment, user preferences, and attention to healthy eating.
Differences between Online and Offline Food Environments (RQ2)
@@
O2O food delivery services extend the service range of local restaurants, though their impact on real-world food swamps remains unclear. In this section, we apply statistical learning algorithms to investigate the spatial clustering of offline and online food environments and compare their relative healthiness.
@@
Figure 6: (a) Spatial clustering based on regions’ local restaurants. (b) Spatial clustering based on regions’ accessible O2O food delivery restaurants. (c) Differences in the number of accessible fast food restaurants across clusters. (d) Differences in the fraction of accessible fast food restaurant across clusters.
@@
We cluster regions (Geohash-6 grids) based on the accessibility of local (offline) and O2O food delivery (online) restaurants, with each region represented by a 104-dimensional vector reflecting the number of accessible restaurants across 104 categories. Local restaurants are defined as those located within 1 kilometer of the region. We employ the K-means clustering algorithm (Lloyd 1982) and use the elbow method (Thorndike 1953) to determine that the optimal number of clusters is five. As shown in Figure 6(a) and (b), both offline and online restaurant accessibility exhibit distinct spatial clustering. We compare the number of accessible fast food restaurants across each cluster in Figure 6(c). The core areas of the city (clusters 1 and 2) have the highest availability of O2O food delivery, with cluster 1, representing two key business districts (Zhongguancun and Guomao), having the highest access.
@@
The number of accessible restaurants online is approximately three times greater than that of offline restaurants, highlighting the pivotal role of O2O food delivery services in expanding access to convenience food options. Figure 6(d) shows the fraction of fast food restaurants ($\phi$) in each cluster for both offline and online food environments. In regions near the city center, the proportion of fast food restaurants is higher. Notably, the proportion of fast food in the online food environment is slightly lower than in the offline environment (with regional averages of 21.2% and 22.1%, respectively), though the online food environment still contains a substantial share of fast food. Moreover, by comparing the red and orange clusters in Figure 6(a) and (b), we observe that O2O food delivery has expanded the size of clusters with the highest fast food count and proportion. This indicates that O2O delivery services may further exacerbate real-world food swamps, potentially creating a “cyber food swamp”.
@@
Figure 7: The average choices on fast food orders of regions within different levels of offline (a) and online (b) food environment $\phi$, as represented by the fraction of accessible fast food restaurants. Error bars indicate 95% confidence intervals.

### v3

@@ Abstract
Online-to-offline (O2O) food delivery platforms have greatly expanded urban residents’ access to a wide range of food options by allowing convenient ordering from distant food outlets. However, concerns persist regarding the nutritional quality of delivered food, particularly as the impact of O2O food delivery platforms on users’ healthy food remains unclear. This study leverages large-scale empirical data from a leading O2O delivery platform to comprehensively analyze online food choice behaviors and how they are influenced by the online exposure to fast food restaurants, i.e., online food environment. Our analyses reveal significant variations in food preferences across demographic groups and city sizes, where male, low-income, and younger users are more likely to order fast food via O2O platforms. Besides, we also perform a comparative analysis on the food exposure differences in offline and online environments, confirming that the extended service ranges of O2O platforms can create larger “cyber food swamps”. Furthermore, regression analysis highlights that a higher ratio of fast food orders is associated with “cyber food swamps”, areas characterized by a higher proportion of accessible fast food restaurants. A 10% increase in this proportion raises the probability of ordering fast food by 22.0%. Moreover, a quasi-natural experiment substantiates the long-term causal effect of online food environment changes on healthy food choices. These findings underscore the need for O2O food delivery platforms to address the health implications of online food choice exposure, offering critical insights for stakeholders aiming to improve dietary health among urban populations.
Code & Datasethttps://github.com/tsinghua-fib-lab/CyberFoodSwamp
@@ Introduction
Although the convenience of O2O food delivery is undeniable, concerns persist regarding the nutritional quality of delivered food, with numerous criticisms pointing to its generally low nutritional value (Dai, Wu, and Hu 2022). Drawing on the concept of “food swamps”, which describes areas oversaturated with unhealthy food outlets in the physical world (Stowers, Schwartz, and Brownell 2017), O2O food delivery platforms have the potential to expand food access and create virtual food swamps that shape consumer choices towards unhealthy restaurants, leading to adverse health outcomes (Stephens, Miller, and Militello 2020). Therefore, it is essential to assess how online food environments on these platforms influence users’ healthy food choices.
Existing studies primarily used surveys and field studies to evaluate O2O food delivery platforms from the perspective of healthiness, including the nutritional value of delivered food (Brar and Minaker 2021; Dai, Wu, and Hu 2022), consumer perceptions of healthy food availability (Dai, Wu, and Hu 2022; Eu and Sameeha 2021; Osaili et al. 2023), and healthy food choices (Osaili et al. 2023; Giacomini et al. 2024). These studies often suffered from limitations in the scale of research samples and lack a direct link between online food environment and consumers’ real-world food choices. Furthermore, while significant efforts have been made to examine the impact of traditional dining environments on food choices (Feng et al. 2010; Althoff et al. 2022; García Bulle Bueno et al. 2024), their findings may not directly correspond to the effects of online food environments. This is particularly relevant as O2O food delivery significantly expands the spatial range of food options, highlighting the need to compare offline and online food swamps.
@@
To answer the above questions, we conduct comprehensive data-driven analyses and derive insightful findings based on O2O food delivery restaurant and consumption data at both city and individual levels, collected from one of the largest O2O food delivery platforms in China. First, we focus on the macro-level health discrepancies within the O2O food delivery industry. We uncover a scaling law that governs the relation between O2O food delivery consumptions and city sizes, showing larger cities have significantly higher proportions of fast food restaurants and placed delivery orders, which may be attributed to the faster pace of social life in these cities (Bettencourt et al. 2007). Healthiness of food choices is also correlated with demographic, with males, individuals with lower incomes, and younger people exhibiting a significantly higher likelihood of ordering fast food online. This trend, derived from large-scale empirical data, mirrors previous findings on food destinations based on small-sample surveys (Kerr et al. 2012). Second, we compare the distributional patterns between offline and online food environments. The number of accessible O2O restaurants is significantly higher than that of local offline restaurants, while the healthiness of the online food environment is slightly better than that of the offline environment. However, the extended service range of O2O platforms expands the spatial coverage of food swamps. The online food environment shows a stronger correlation with healthy food choices on O2O food delivery platforms. Finally, following the analytic framework proposed by García Bulle Bueno et al. (García Bulle Bueno et al. 2024), we examine the relationship between online food environments and healthy food choices. Logistic regressions reveal that a higher proportion of fast food restaurants in a user’s location is associated with a higher likelihood of ordering fast food. Notably, low-income individuals and younger users are more affected by this environment, underscoring the heterogeneous susceptibility of users on online platforms (Li et al. 2022; Sukiennik, Gao, and Li 2024). A quasi-natural experiment based on users who permanently change their ordering context further substantiates the gradual, negative causal impact of online food environments dominated by unhealthy fast food options on healthy dining habits, highlighting the urgent need for regulations addressing “cyber food swamps”.
Our main contributions can be summarized as follows.
@@
$\bullet$ We leverage empirical O2O food delivery consumption data to comprehensively understand the healthiness of the online food environment and its impact on healthy food choices, highlighting the significant potential for exploring the interplay between the web and society using data-driven methods.
@@
$\bullet$ We demonstrate macro-level health discrepancy within the online food environment by uncovering city-level scaling laws of O2O food delivery adoption and identifying disparities in healthy food choices across demographic groups.
$\bullet$ We compare the healthiness of offline and online food environments and confirm the amplifying effect of O2O food delivery on real-world food swamps.
@@
$\bullet$ We reveal the effects of online food environments on healthy food choice behaviors through a quasi-intervention experiment utilizing a detailed, individual order-level dataset, indicating the negative health impacts of “cyber food swamps”.
@@
Many previous studies and reports have substantiated the rapid growth of O2O food delivery in global cities, highlighting it as a convenient and viable solution to address limited access to healthy food options (Dillahunt, Simioni, and Xu 2019; Yeo, Goh, and Rezaei 2017). During the COVID-19 pandemic, social distancing measures made O2O food delivery a primary means of obtaining food, further accelerating the industry’s growth (Dsouza and Sharma 2021; Li, Mirosa, and Bremer 2020; Hong et al. 2021). At the same time, the nutritional and health implications of O2O food delivery have attracted increasing attention from researchers worldwide. As for users’ perceptions of delivered food, surveys conducted in Jordan (Osaili et al. 2023), Malaysia (Eu and Sameeha 2021), and China (Dai, Wu, and Hu 2022) found that most participants believe food delivered through O2O food delivery is generally less healthy than food served in restaurants. However, many of these participants still report a high reliance on such services. Several studies have evaluated the healthiness of O2O food delivery meals through sampling. An analysis of food delivery menus in Canada revealed that the majority of menus scored below the standards set by the Healthy Eating Index (Brar and Minaker 2021). Similarly, a study on the nutritional quality of food sold on Chinese O2O food delivery platforms found that most popular items have low nutritional value (Dai, Wu, and Hu 2022) – younger consumers, in particular, tend to pay less attention to the nutritional content of delivered meals, which can lead to health issues such as elevated cholesterol and obesity with long-term consumption. In Italy, an analysis of differences in willingness to use food delivery services among users found that individuals with lower health literacy are more dependent on O2O food delivery (Giacomini et al. 2024). Most current studies are limited to small-scale data collection through surveys or sampling, without assessing the overall health preferences of O2O food delivery users across a broader population. In our study, we leverage a large-scale empirical dataset covering the consumption behaviors of hundreds of thousands of users to comprehensively analyze their online food environments and healthy food choices.
@@
Digital data collected from online services, including activity tracking, GPS traces, and social media posts, have been widely linked to the health outcomes of urban residents in existing research (Lin et al. 2020; Zhang et al. 2021a, b; Xu et al. 2025). Large-scale physical mobility data collected via smartphone applications has been used to reveal the discrepancy in urban residents’ health status (Chen et al. 2022; Zhang et al. 2024, 2025a). For example, such discrepancy significantly decreases as city walkability improves (Althoff et al. 2017). A study based on mobile GPS data found that while large cities offer more recreational and social options, they tend to direct individuals to locations frequented by people of similar socioeconomic status, increasing the segregation of social activities and leading to new public health risks, such as lack of physical activity and imbalanced diets (Nilforoshan et al. 2023; Chen et al. 2023; Fan et al. 2025). Social media data, spanning multiple modalities, are also emerging as promising proxies for public health indicators. Posts on X (formerly Twitter) have been used to predict county-level public health metrics such as obesity and diabetes rates (Culotta 2014; Abbar, Mejova, and Weber 2015). Food-related photos on Instagram, combined with location information, can accurately reflect offline food deserts (De Choudhury, Sharma, and Kiciman 2016). YouTube videos have been applied to track dietary changes during the COVID-19 pandemic (Mejova and Manikonda 2023). Additionally, online recipe platforms are being used to nudge healthier food adoption by promoting specific healthy recipes (Jesse, Jannach, and Gula 2021; Chelmis and Gergin 2023). Similarly, data generated by O2O food delivery platforms hold the potential to reflect the health-related information of urban residents.
@@
In this study, we mainly use three categories of data, i.e., O2O food delivery restaurant data, O2O food delivery consumption order data, and demographic data.
@@
We collect information on O2O food delivery restaurants from one of China’s largest O2O food delivery platforms. For each restaurant, the platform provides its name, category, and service range. Each restaurant is assigned to a single, specific category. To assess the healthiness of the restaurants, we classify those within the “fast food and snacks”, “western fast food”, and “fried food” categories, as these are often regarded as having low nutritional value and may contribute to health issues such as obesity (Hu et al. 2016). All other categories are considered non-fast food restaurants. On average, 18.77% of restaurants are categorized as fast food restaurant.
@@
The service range of each O2O food delivery restaurant indicates the urban areas that can access its delivery services. As depicted in Figure 1, the restaurant marked in blue has a broader service range than the one marked in pink. Utilizing this data, we can identify O2O food delivery restaurants accessible to each urban area. We divide the city of Beijing into Geohash-6 “regions” (approximately 1.2$\times$0.6 kilometers), represented as rectangles in Figure 2(a). For each region $c$, the “unhealthiness” of its corresponding online food environment $\phi(c)$ is defined as the ratio of the number of accessible fast food restaurants to the total number of accessible restaurants. Figure 2(a) illustrates the spatial distribution of $\phi(c)$ across 4,921 regions in Beijing.
@@
The city-level dataset records the number of daily O2O food delivery orders, including fast food orders, across 283 Chinese cities in 2018 and 2024, respectively. For the individual-level dataset, we randomly sample 850,000 platform users in Beijing and collect their order records from January 2017 to May 2018. Additionally, we gathered order records from 20,000 users in Beijing, 10,000 users in Chengdu (a second-tier city in southwestern China), and 10,000 users in Xiamen (a third-tier city in southeastern China) for the period from January 2023 to December 2024. Each entry includes the user ID, order time, user location (at the Geohash-6 region level), expenditure amount, and the corresponding O2O food delivery restaurant. For the largest Beijing dataset from 2018, each user placed 50.3 orders on average. Based on these detailed order records, we calculate the fast food preference $\psi$ of each user, defined as the ratio of fast food orders to the total number of orders. Figure 2(b) illustrates the distribution of $\psi$ on the sampled population. On average, 27.98% of O2O food delivery orders are placed at fast food restaurants.
@@
To establish a connection between healthy food choices and user attributes, we further obtain demographic data from the platform, which includes gender, income level, and age, as inferred by the platform’s machine learning algorithms from user-provided profiles and user behaviors. We filter out users whose demographic profiles have a confidence level below 95%. To ensure user privacy, income levels are categorized into three groups: “high-income”, “medium-income” and “low-income”. Similarly, age groups are segmented into “below 25”, “25 to 40”, and “over 40”. The demographic distributions of the sampled population are depicted in Figure 3.
@@
We first demonstrate the macroscopic city-level health discrepancy of O2O food delivery. Urban services and economic activities often follow a scaling law, expressed as $Y_{i}=Y_{0}N_{i}^{\alpha}$ (Bettencourt 2021; Xu et al. 2021; Ribeiro and Rybski 2023), where $Y_{i}$ represents a measure of socioeconomic activity or resources, and $N_{i}$ denotes the city $i$’s population. An exponent $\alpha<1$ indicates sublinear growth of $Y$ relative to the city population, typically associated with urban infrastructure such as gas stations or road length, where the per-capita infrastructure decreases in larger cities (Um et al. 2009). In contrast, socioeconomic factors like GDP or crime rates often exhibit $\alpha>1$, signifying superlinear growth, where per-capita values increase more rapidly in larger cities due to intensified social interactions (Bettencourt 2013; Succar and Porfiri 2024).
Figure 4: Scaling laws of O2O food delivery restaurants (a), fast food restaurants (b), O2O food delivery orders (c), and fast food orders (d) in Chinese cities in 2018. The axes are in logarithmic scales.
We collect city population from China City Statistical Yearbook [^1] https://www.stats.gov.cn/sj/ndsj/2019/indexch.htm and analyze the scaling law of O2O food delivery. As an innovative form of urban services, O2O food delivery in Chinese cities also exhibits scaling behavior. The fitted lines in Figure 4(a) and (c) indicate that the scaling exponent $\alpha$ for the number of O2O food delivery restaurants and the number of O2O food delivery orders are 0.86$\pm$0.07 and 1.17$\pm$0.10, respectively. This corresponds to the typical sublinear and superlinear growth patterns of urban infrastructure and economic activities, with $\beta=\frac{5}{6}$ and $\frac{7}{6}$, reflecting higher infrastructure utilization efficiency in large cities (Bettencourt 2013). Similarly, these patterns are also observed in the 2024 dataset, as depicted in Figure 12.
Furthermore, the scaling exponents for fast food restaurants and orders in 2018 are significantly higher than those for all food categories ($P<0.05$, two-sided t-test), as shown in Figure 4(b) and (d). Specifically, doubling the city population increases the proportion of fast food restaurants by 4.0% and the proportion of fast food orders by 5.4%. This difference suggests a health discrepancy across cities, with larger cities demonstrating a greater reliance on fast food restaurants through O2O food delivery (Bettencourt et al. 2007). However, this trend is not observed in 2024 (Figure 12(b) and (d)), where the scaling coefficients for all food categories and fast food restaurants show no significant differences, indicating a shift toward healthier dining habits among metropolitan residents post-pandemic.
@@
The observed health discrepancies across cities of different sizes may be related to variations in the composition of urban populations. To test this assumption, we further investigate the relationships between user demographics and their behaviors on O2O food delivery platforms. We focus on three demographic attributes – gender, income level, and age – and compare the average proportion of fast food orders and the average order price for each group. The values and corresponding 95% confidence intervals are shown in Figure 5.
@@
Figure 5: The average fast food order ratio and order price across different gender (a), income level (b), and age (c) groups in Beijing. Error bars represent 95% confidence intervals.
@@
The proportion of fast food orders among male users (33.0%) is higher than that of female users (25.0%) in 2018, while the average price of food delivery orders is nearly identical for males (58.4 CNY) and females (58.6 CNY). This indicates that male users have a significant stronger preference for fast food delivery compared to female users ($P<0.001$, two-sided t-test). In terms of income level, the average order price increases with higher income, while the proportion of fast food orders decreases. Users of highest income level only have 26.0% fast food orders, while the ratio of the lowest income level reaches 32.0%. This suggests that higher-income groups tend to opt for more expensive, healthy food delivery options, whereas lower-income groups prefer cheaper, less healthy food options ($P<0.001$, two-sided t-test). Similarly, the average price of food delivery orders rises with age, while the proportion of fast food orders declines from 34.8% among users under 25 years to 24.8% for users over 40 years, indicating that younger people are more inclined to choose fast food, while older individuals tend to prefer higher-priced alternatives ($P<0.001$, two-sided t-test).
@@
This overall health discrepancy remains robust across three Chinese cities in 2024, as shown in Figure 13. This consistent trend highlights the heterogeneous healthy food choices of O2O food delivery platform users, which may be influenced by factors such as service pricing (Zhang et al. 2025b), the online food environment, user preferences, and attention to healthy eating.
Differences Between Offline and Online Food Environments (RQ2)
@@
O2O food delivery services extend the service range of local restaurants, though their impact on real-world food swamps remains unclear. In this section, we apply statistical learning algorithms to investigate the spatial clustering of offline and online food environments and compare their unhealthiness.
@@
Figure 6: The comparison of offline and online food swamps. (a) Spatial clustering based on regions’ local restaurants. (b) Spatial clustering based on regions’ accessible O2O food delivery restaurants. (c) Differences in the number of accessible fast food restaurants across clusters. (d) Differences in the proportion of accessible fast food restaurant across clusters.
@@
We cluster regions (Geohash-6 grids) based on the accessibility of local (offline) and O2O food delivery (online) restaurants, with each region represented by a 104-dimensional vector reflecting the number of accessible restaurants across 104 categories. Local restaurants are defined as those located within 1 kilometer of the region. We employ the K-means clustering algorithm (Lloyd 1982) and use the elbow method (Thorndike 1953) to determine that the optimal number of clusters is 5. As shown in Figure 6(a) and (b), both offline and online restaurant accessibility exhibit distinct spatial clustering. We compare the number of accessible fast food restaurants across each cluster in Figure 6(c). The core areas of the city (clusters 1 and 2) have the highest availability of O2O food delivery, with cluster 1, representing two key business districts (Zhongguancun and Guomao), having the highest access.
@@
The number of accessible online restaurants is approximately three times greater than that of offline restaurants, highlighting the pivotal role of O2O food delivery services in expanding access to convenience food options. Figure 6(d) shows the proportion of fast food restaurants ($\phi$) in each cluster for both offline and online food environments. In regions near the city center, the proportion of fast food restaurants is higher. Notably, the proportion of fast food in the online food environment is slightly lower than in the offline environment (with regional averages of 21.2% and 22.1%, respectively), though the online food environment still contains a substantial proportion of fast food. Moreover, by comparing the red and orange clusters in Figure 6(a) and (b), we observe that O2O food delivery has expanded the size of clusters with the highest fast food count and proportion. This indicates that O2O delivery services may further exacerbate real-world food swamps, potentially creating a “cyber food swamp”.
@@
Figure 7: The average choices on fast food orders of regions within different levels of offline (a) and online (b) food environment $\phi$, as represented by the proportion of accessible fast food restaurants. Error bars indicate 95% confidence intervals.